An auto mechanic has 600g of battery acid solution that is 60% acid. he must add water to this solution to dilute it so thst it is only 45% acid. How much water should he add?.

Answer by Theo(13342)

.6 * 600 + 0y = .45 * (600 + y)
simplify to get:
360 = 270 + .45y
solve for y to get:
y = (360 - 270) / .45 = 200
that's your solution.
200 grams of water needs to be added to 60% acid solution to make 45% acid solution.